# TAG-WM: Tamper-Aware Generative Image Watermarking via Diffusion Inversion Sensitivity

### v3

@@ 1 Introduction
The rise of AI-generated content (AIGC) has garnered significant attention across various fields, creating substantial commercial value. Particularly in visual content generation, the advent of diffusion models [3, 4, 21] has sparked the emergence of numerous image generation and manipulation applications [29], enabling individuals across industries to easily and efficiently perform customized generation or editing of images with high quality. However, this technological accessibility inevitably leads to uncertainties in the provenance and authenticity of images in the AIGC era, resulting in copyright risks and disinformation threats. For instance, users might use image generators to create similar graphics to copyrighted work and falsely claim ownership. Furthermore, with image generators, malicious actors could easily create or manipulate lifelike images to spread false information about non-existent events.
@@
As a common technique for integrity verification and source tracing, digital image watermarking has gained increasing prominence in AIGC. Existing image watermarking methods were primarily designed by traditional image processing algorithms [14, 1] or deep neural networks [13, 8], both of which embed robust watermarks through post-processing cover images. However, such post-processing embedding approaches inevitably introduce visual artifacts. Considering that the quality of generated images is the goal pursued by image generation models at a high computational cost, such watermarking methods contradict the application scenarios of generative models. With the rise of diffusion models, new watermarking paradigms have emerged to protect AI-generated content. Recent advancements focus on model fine-tuning [6, 9, 26] and latent space-based watermarks [23], watermarking directly during the generative process itself. Such watermarking methods embedded in the generative process can better integrate watermark information with image content while significantly reducing the impact of watermark embedding on the visual quality of generated images. The Gaussian Shading (GS) proposed by Yang et al. [27] takes a significant step forward, achieving provably visual quality lossless watermarking for generated images. By employing distribution-preserving sampling to map watermarks into latent space representations, GS ensures the watermarked latent becomes statistically indistinguishable from the original one, thereby maintaining the visual quality equivalent to the original generative images.
@@
Meanwhile, the emergence of generative models like ControlNet [29], which enable high-quality image modifications, has exacerbated the threat of malicious tampering, imposing stricter requirements for watermark robustness and tamper localization capabilities. Unlike common image distortions, malicious alterations fundamentally alter pixel values. Unfortunately, the DDIM inversion process demonstrates sensitivity to pixel modifications, rendering inversion-based generative watermarking methods like GS less robust against image manipulation distortions (e.g., cropping, tampering). On the other hand, as generative model-based image manipulations become increasingly indistinguishable, applying watermarks to achieve more generalizable proactive tamper localization has emerged as a novel and practical demand in watermark functionality. For instance, MaLP [2] achieves tamper detection and pixel-level localization through learned template embedding. EditGuard [32] simultaneously embeds both identification watermarks and tampering localization watermarks into images, decoding them in parallel to accomplish dual objectives of copyright verification and tamper localization.
@@ 2 Related Work
Diffusion-based Image Generation The explosive growth of diffusion models has revolutionized image synthesis, with frameworks like DDPM [12] enabling high-fidelity generation, DDIM [22] enabling fast deterministic sampling via non-Markov processes, LDMs [21] compressing data into latent space for efficiency, text-guided models like Stable Diffusion and Imagen [3] leveraging large-scale pretraining for diverse generations, while ControlNet [29] facilitates controllable creation and manipulation. These advancements have made diffusion models the standard for generative tasks, yet their very success intensifies the urgency of addressing copyright risks in generated content.
@@
Watermarking for Generative Models While post-processing methods [35, 33, 25] suffer from capacity-quality trade-offs, Tree-Ring [23] proposed in-generation approaches for popular DMs, which bases on the non-Markov process and deterministic of DDIM inversion [22]. Gaussian Shading [27] uses it and achieves provable quality-lossless multi-bit watermarking. However, their reliance on noise patterns makes them vulnerable to image manipulations — a critical limitation given the prevalence of image editing tools. This motivates our tamper-aware watermarking that inherits in-generation advantages while resisting post-hoc manipulations.
@@
Tampering Localization Passive methods detect specific manipulation traces [24] or general artifacts [7]. As diffusion-based manipulation [19, 29] increasingly produces undetectable forgeries, passive detectors face diminishing returns. Proactive solutions like EditGuard [32] circumvent this issue by employing watermarking techniques, but they suffer from generation-time overheads, image quality degradation, and limited robustness against out-of-distribution (OOD) samples (e.g., untrained degraded inputs). Our approach addresses these limitations by embedding localization watermarks directly into the diffusion pipeline, ensuring seamless integration without compromising efficiency or fidelity.
@@ 3.1 Implementation Details
Benckmark Dataset. During inference, we employ the prompt from Stable-Diffusion-Prompt [^1] https://huggingface.co/datasets/Gustavosta/Stable-Diffusion-Prompts, with a guidance scale of 7.5. We sample 50 steps using DPMSolver [18]. Considering that users tend to propagate the generated images without retaining the corresponding prompts, we use an empty prompt for inversion, with a scale of 1. We perform 50 steps of inversion using DDIM inversion [22]. Through that, we get our test set includes 1,000 images, which are kept completely isolated from the training and validation sets. All degraded data samples were randomly selected from the predefined degradation types and intensity levels specified in Appendix A.
@@ 3.2 Watermark Performance
Baseline Methods. We select seven baseline methods: three officially used by SD, namely DwtDct [5] and DwtDctSvd [5]; two post-processing-based methods RivaGAN [28] and EditGuard [32], a fine tuning-based method called Stable Signature [9], a latent representation-based method called Tree-Ring [23], and GS [27].
Evaluation Metrics. To measure the performance of watermarking methods, we calculated the bit accuracy (Bit Acc). To measure the bias in model performance, we computed the CLIP-Score [20] for 10 batches of watermarked images and performed a t-test on the mean CLIP-Score compared to that of watermark-free images. In prior works, the incorporation of watermark embedding modules inevitably results in a decline in model performance, therefore typically evaluated using the Peak Signal-to-Noise Ratio (PSNR) and Fréchet Inception Distance (FID) [11], which are unnecessary to our method. To simulate detecting and tracing scenarios, we also calculated the detecting true positive rate (D-TPR) and the tracing true positive rate (T-TPR), by fixing the false positive rate (FPR) and the number of users at $10^{-6}$ and $10^{6}$, separately [27].

### v4

@@ 1 Introduction
The rise of AI-generated content (AIGC) has garnered significant attention across various fields, creating substantial commercial value. Particularly in visual content generation, the advent of diffusion models [DMs-Imagen3, DMs-DALLE3, DMs-LDMs] has sparked the emergence of numerous image generation and manipulation applications [DMs-ControlNet], enabling individuals across industries to easily and efficiently perform customized generation or editing of images with high quality. However, this technological accessibility inevitably leads to uncertainties in the provenance and authenticity of images in the AIGC era, resulting in copyright risks and disinformation threats. For instance, users might use image generators to create similar graphics to copyrighted work and falsely claim ownership. Furthermore, with image generators, malicious actors could easily create or manipulate lifelike images to spread false information about non-existent events.
@@
As a common technique for integrity verification and source tracing, digital image watermarking has gained increasing prominence in AIGC. Existing image watermarking methods were primarily designed by traditional image processing algorithms [old_wm-wavelet, old_wm-combined_DWT_DCT] or deep neural networks [postprocessing_watermarking-MBRS, postprocessing_watermarking-PIMoG], both of which embed robust watermarks through post-processing cover images. However, such post-processing embedding approaches inevitably introduce visual artifacts. Considering that the quality of generated images is the goal pursued by image generation models at a high computational cost, such watermarking methods contradict the application scenarios of generative models. With the rise of diffusion models, new watermarking paradigms have emerged to protect AI-generated content. Recent advancements focus on model fine-tuning [fine_tuning_based_watermarking-Diffusionshield, fine_tuning_based_watermarking-StableSignature, fine_tuning_based_watermarking-flexible] and latent space-based watermarks [latent_representation_based_watermarking-Treering], watermarking directly during the generative process itself. Such watermarking methods embedded in the generative process can better integrate watermark information with image content while significantly reducing the impact of watermark embedding on the visual quality of generated images. The Gaussian Shading (GS) proposed by Yang et al. [gaussianshading] takes a significant step forward, achieving provably visual quality lossless watermarking for generated images. By employing distribution-preserving sampling to map watermarks into latent space representations, GS ensures the watermarked latent becomes statistically indistinguishable from the original one, thereby maintaining the visual quality equivalent to the original generative images.
@@
Meanwhile, the emergence of generative models like ControlNet [DMs-ControlNet], which enable high-quality image modifications, has exacerbated the threat of malicious tampering, imposing stricter requirements for watermark robustness and tamper localization capabilities. Unlike common image distortions, malicious alterations fundamentally alter pixel values. Unfortunately, the DDIM inversion process demonstrates sensitivity to pixel modifications (Sec. 3.2), rendering inversion-based generative watermarking methods like GS less robust against image manipulation distortions (e.g., cropping, tampering). On the other hand, as generative model-based image manipulations become increasingly indistinguishable, applying watermarks to achieve more generalizable proactive tamper localization has emerged as a novel and practical demand in watermark functionality. For instance, MaLP [tl-proactive-MaLP] achieves tamper detection and pixel-level localization through learned template embedding. EditGuard [wm_and_tl-Editguard] simultaneously embeds both identification watermarks and tampering localization watermarks into images, decoding them in parallel to accomplish dual objectives of copyright verification and tamper localization.
@@ 2 Related Work
Diffusion-based Image Generation The explosive growth of diffusion models has revolutionized image synthesis, with frameworks like DDPM [DMs-DDPMs] enabling high-fidelity generation, DDIM [DMs-scheduler-DDIM] enabling fast deterministic sampling via non-Markov processes, LDMs [DMs-LDMs] compressing data into latent space for efficiency, text-guided models like Stable Diffusion and Imagen [DMs-Imagen3] leveraging large-scale pretraining for diverse generations, while ControlNet [DMs-ControlNet] facilitates controllable creation and manipulation. These advancements have made diffusion models the standard for generative tasks, yet their very success intensifies the urgency of addressing copyright risks in generated content.
@@
Watermarking for Generative Models While post-processing methods [postprocessing_watermarking-HiDDeN, postprocessing_watermarking-REVMark, postprocessing_watermarking-SepMark] suffer from capacity-quality trade-offs, Tree-Ring [latent_representation_based_watermarking-Treering] proposed in-generation approaches for popular DMs, which bases on the non-Markov process and deterministic of DDIM inversion [DMs-scheduler-DDIM]. Gaussian Shading [gaussianshading] uses it and achieves provable quality-lossless multi-bit watermarking. However, their reliance on noise patterns makes them vulnerable to image manipulations — a critical limitation given the prevalence of image editing tools. This motivates our tamper-aware watermarking that inherits in-generation advantages while resisting post-hoc manipulations.
@@
Tampering Localization Passive methods detect specific manipulation traces [tl-passive-specific-OSN] or general artifacts [tl-passive-universal-MVSSNet]. As diffusion-based manipulation [ImageInpainting-SDXL, DMs-ControlNet] increasingly produces undetectable forgeries, passive detectors face diminishing returns. Proactive solutions like EditGuard [wm_and_tl-Editguard] circumvent this issue by employing watermarking techniques, but they suffer from generation-time overheads, image quality degradation, and limited robustness against out-of-distribution (OOD) samples (e.g., untrained degraded inputs). Our approach addresses these limitations by embedding localization watermarks directly into the diffusion pipeline, ensuring seamless integration without compromising efficiency or fidelity.
Figure 1: The proposed TAG-WM framework. It embeds copyright $W_{cop}$ and localization watermarks $W_{loc}$ through dual-watermark joint sampling strategy. By analyzing dense variation regions of $W_{loc}$, it enables tamper localization while improving watermark decoding accuracy using tampering insights.
3 Preliminaries
3.1 Denoising Diffusion Implicit Model
The Denoising Diffusion Implicit Model (DDIM) accelerates diffusion-based generation by defining a non-Markovian process. It enables deterministic sampling through a modified reverse process. The generation process can be formulated as:
$x_{t-1}=\sqrt{\frac{\bar{\alpha}_{t-1}}{\bar{\alpha}_{t}}}x_{t}+(\sqrt{1-\bar{\alpha}_{t-1}}-\sqrt{\frac{\bar{\alpha}_{t-1}(1-\bar{\alpha}_{t})}{\bar{\alpha}_{t}}})\epsilon_{\theta}^{(t)}(x_{t})$
Suppose $\epsilon_{\theta}^{(t)}(x_{t})\approx\epsilon_{\theta}^{(t-1)}(x_{t-1})$, then the inversion process is as follows:
$x_{t}\approx\sqrt{\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}}}x_{t-1}+(\sqrt{1-\bar{\alpha}_{t}}-\sqrt{\frac{\bar{\alpha}_{t}(1-\bar{\alpha}_{t-1})}{\bar{\alpha}_{t-1}}})\epsilon_{\theta}^{(t-1)}(x_{t-1})$
which we used to recover watermarked pseudo-noise.
3.2 Spatial Mapping Properties
According to the mapping relationship Eq. (2) between the noise $x_{t}$ and the image $x_{0}$ in the DDIM framework, it is evident that any modification to $x_{0}$ in pixel space will propagate to the corresponding positions in $x_{t}$. This property is preserved to some extent in the VAE-based latent-diffusion architecture [DMs-LDMs], as the VAE retains spatial structure while accelerating inference. Consequently, spatial perturbations in the image domain induce localized changes in the latent space during inversion.
4 Method
As illustrated in Fig. 1, the proposed TAG-WM framework operates through four stages. During watermark embedding, our method encodes both copyright watermark $W_{cop}$ and localization watermark $W_{loc}$ via a dual-watermark joint sampling strategy integrated into the diffusion model generation process. For watermark extraction, the framework first reconstructs latent watermark components through inverse sampling. Subsequently, it performs tampering localization by analyzing dense variation regions in $W_{loc}$, while simultaneously leveraging this tampering information to enhance the decoding accuracy of copyright messages.
4.1 Dual Watermark Embedding
Watermark Initialization
The watermark generation process involves creating both copyright watermark $W_{cop}$ and localization watermark $W_{loc}$. For $W_{cop}$, the copyright messages $m\in\{0,1\}^{L}$ are adaptively expanded to match the dimensional space of the latent representations in generation models. For latent variables with dimensions $C\times H\times W$, the expansion process operates as follows:
$\hat{m}=\Bigl[\,\overbrace{m,\dots,m}^{N},\,m_{1:R}\,\Bigr]$ (3)
where $N=\lfloor D/L\rfloor$ represents full replication count and $R=D\bmod L$ corresponds to residual bits, $D=C\times H\times W$. Then, the expanded message $\hat{m}$ is encrypted via ChaCha20 [chacha20] with key $k$ to obtain $W_{cop}$ with a multiple binary uniform distribution:
$W_{cop}=\mathrm{Reshape}\left(\mathrm{ChaCha20}(\hat{m},k),(C,H,W)\right)$ (4)
For the localization watermark $W_{loc}$, a deterministic pseudo-random process is employed. Using a fixed seed $s$, a pseudo-random generator $G$ is adopted to generate $W_{loc}$:
$W_{loc}=\mathrm{Reshape}\left(G(s,\theta),\ (C,H,W)\right)$ (5)
where each element follows $\mathcal{B}(1-\theta)$. $\mathcal{B}$ donates the Bernoulli distribution, and the $\theta$ is the probability of zeros.
Dual-Mark Joint Sampling
To embed dual watermarks while preserving visual quality, we propose the Dual-Mark Joint Sampling (DMJS) algorithm, which generates dual-watermarked pseudo-Gaussian noise $Z_{T}^{w}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ according to $W_{cop}$ and $W_{loc}$. Each element $z$ in $Z_{T}^{w}$ corresponds to the bit combination $(w_{c},w_{l})\in\{(0,0),(0,1),(1,0),(1,1)\}$ of $W_{cop}$ and $W_{loc}$, with its distribution expressed as:
$p(z)=\sum_{(w_{c},w_{l})\in\{0,1\}^{2}}P(w_{c},w_{l})\cdot p(z|w_{c},w_{l})$ (6)
Then, the real number set $\mathbb{R}$ can be partitioned into four continuous subsets for watermark embedding: $(-\infty,a)$, $[a,0)$, $[0,b)$, $b,+\infty)$. A straightforward embedding idea is to create a direct mapping of these four intervals to the bit combination $(0,0),(0,1),(1,0),(1,1)$. However, to mitigate boundary errors from DDIM inversion and noise prediction, we reduce the intervals to three: $(-\infty,a)$, $[a,b)$, $b,+\infty)$ by reordering the embedding sequence to $(0,0),(1,0),(1,1),(1,0)$. This retains two intervals for $W_{cop}$ and three for $W_{loc}$; see Fig. 2.
(a) Four-interval strategy
(b) Three-interval strategy
Figure 2: Embedding strategies for bit pairs $(w_{c},w_{l})$.
Both types of strategies can ensure $Z_{T}^{w}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, thereby preserving the generation quality without degradation. The detailed proof is as follows. The conditional density for $Z_{T}^{w}$ is:
$p(z|w_{c},w_{l})=\frac{\phi(z)}{\mid\Phi(l_{w_{c\&l}})-\Phi(u_{w_{c\&l}})\mid}\cdot\mathbf{1}_{[l_{w_{c\&l}},u_{w_{c\&l}})}(z)$ (7)
where $\phi(\cdot)$ and $\Phi(\cdot)$ are the PDF and CDF of $\mathcal{N}(0,1)$, respectively. $\mathbf{1}_{[a,b)}(\cdot)$ denotes the indicator function, which equals 1 if $z$ lies in the interval $[a,b)$ and 0 otherwise. The interval lower boundary $l_{w_{c\&l}}$ and the interval upper boundary $u_{w_{c\&l}}$ are strategy-dependent as shown in 1, in which $a$ and $b$ are defined as:
$a=\Phi^{-1}(\tfrac{\theta}{2}),b=\begin{cases}\Phi^{-1}(\tfrac{1}{2}+\tfrac{\theta}{2})&\text{(four intervals)}\\ \Phi^{-1}(1-\tfrac{\theta}{2})&\text{(three intervals)}\end{cases}$ (8)
So the probability distribution of $z$ can be denoted as:
$\begin{aligned} p(z)&=\sum_{(w_{c},w_{l})\in\{0,1\}^{2}}\theta^{1-w_{l}}(1-\theta)^{w_{l}}\cdot\tfrac{1}{2}\cdot p(z|w_{c},w_{l})\\ &=\phi(z)\end{aligned}$
confirming $Z_{T}^{w}$ maintains the standard normal distribution required by diffusion models.
Finally, we employ Ordinary Differential Equations (ODEs) [DMs-scheduler-DDIM] to denoise $Z_{T}^{w}$ to obtain the denoised watermarked latent $Z_{0}^{w}$, which is subsequently decoded by the VAE to generate the final watermarked image $X^{w}$.
4.2 Watermark Latent Reconstruction
In the extraction, we first reconstruct watermark latents from a tampered and degraded image ($\widetilde{X}^{w}$). To restore the sampled noise, we first use VAE to encode $\widetilde{X}^{w}$ into latent space, resulting in $\widetilde{Z}_{0}^{w}$. Next, we apply DDIM inversion to obtain the restored watermarked pseudo-Gaussian noise $\widetilde{Z}_{T}^{w}$, which resembles the sampled noise $Z_{T}^{w}$. We then reverse the sampling process according to the boundaries specified in Table 1. Based on the number of intervals chosen and the boundaries associated with each reversed bit, we can reconstruct the watermarks, denoted as $\widetilde{W}_{cop}$ and $\widetilde{W}_{loc}$.
   $(w_{c},w_{l})$    Boundaries $(l_{w_{c\&l}},u_{w_{c\&l}})$
   Four intervals    Three intervals
   $(0,0)$    $(-\infty,a)$    $(-\infty,a)$
   $(0,1)$    $[a,0)$    $[a,0)$
   $(1,0)$    $[0,b)$    $b,+\infty)$
   $(1,1)$    $b,+\infty)    [$0,b)[0,b)
Table 1: Sampling boundaries under different strategies.
4.3 Tampering Localization
The proposed TAG-WM localizes the tampered regions using the reconstructed localization watermark $\widetilde{W}_{loc}$ and the original $W_{loc}$ generated by the shared seed $s$. First, the initial tampering estimation $\widetilde{M}_{tam}^{ini}$ can be calculated through XOR operation between $W_{loc}$ and $\widetilde{W}_{loc}$. To refine the localization result, we propose a dense variation region detector (DVRD), motivated by the observation that local tampering disrupts the deterministic relationship between $W_{loc}$ and $\widetilde{W}_{loc}$, thereby randomizing the bit distribution at latent space positions corresponding to the tampered regions.
This randomization process can be modeled as follows. Let event $A$ denote tampering in the image region corresponding to $\widetilde{w}_{l}\ (\in\widetilde{W}_{loc})$, defined as:
$A=\left\{\widetilde{w}_{l}\,\middle|\,\parbox{173.90271pt}{The image region corresponding to $\widetilde{w}_{l}$ is tampered.}\right\}$
Then, the conditional probabilities are expressed as:
$P(\widetilde{w}_{l}=0\mid A)=\theta,\quad P(\widetilde{w}_{l}=1\mid A)=1-\theta$ (10)
The error probability under tampering becomes:
$P(\widetilde{w}_{l}\neq w_{l}\mid A)=2\theta(1-\theta)$ (11)
which is a concave function that peaks at 0.5 when $\theta=0.5$. Setting $\theta=0.5$, the intrinsic error probability $P(\widetilde{w}_{l}\neq w_{l}\mid\bar{A})$ is experimentally measured as 0.14513. The significant gap between maximum tampering-induced errors (0.5) and intrinsic errors (0.14513) enables threshold-based detection through analysis of dense variation region in $\widetilde{M}_{tam}^{ini}$.
The DVRD algorithm has two implementations: 1) Train-free DVRD computes densities of variation region via multi-scale convolution kernels to balance fine/coarse detection. However, manual kernel optimization proves impractical due to infinite kernel size combinations. 2) Trainable DVRD addresses this limitation using a UNet architecture that automatically learns optimal thresholds and weights across scales through downsampling/upsampling operations and skip connections. By training the detection model, the proposed method can refine coarse tamper localization masks $\widetilde{M}_{tam}^{ini}$ into latent tensor $\widetilde{M}_{tam}^{ref}$ that more accurately aligns with the image tampering regions without manual configuration.
Training methodology and comparative performance analysis between both DVRD variants are detailed in Sec 5.4 and Sec 5.5, respectively. The final output is the refined tampering localization mask $\widetilde{M}_{tam}^{ref}$. To get the binary tampering localization mask $\widetilde{M}_{tam}$ in image space, we decode $\widetilde{M}_{tam}^{ref}$ through the VAE decoder and binarize the decoded output by applying a threshold of 0.
4.4 Tamper-Aware Copyright Message Decoding
Finally, we decode copyright messages from $\widetilde{W}_{cop}$, using the refined tampering mask $\widetilde{M}_{tam}^{ref}$ as a guide to enhance robustness against tampering distortions. Specifically, we first decrypt $\widetilde{W}_{cop}$ with the original encryption key and then apply tamper-aware decoding (TAD). By excluding watermark bits $w_{c}$ identified as compromised in $\widetilde{M}_{tam}^{ref}$, the system aggregates remaining reliable bits through majority voting to reconstruct $\widetilde{m}$. This selective exclusion of tampered regions significantly improves copyright message recovery accuracy under malicious modifications.
@@ 5.1 Implementation Details
Benckmark Dataset. During inference, we employ the prompt from Stable-Diffusion-Prompt [^1] https://huggingface.co/datasets/Gustavosta/Stable-Diffusion-Prompts, with a guidance scale of 7.5. We sample 50 steps using DPMSolver [DMs-scheduler-DPMSolver]. Considering that users tend to propagate the generated images without retaining the corresponding prompts, we use an empty prompt for inversion, with a scale of 1. We perform 50 steps of inversion using DDIM inversion [DMs-scheduler-DDIM]. Through that, we get our test set includes 1,000 images, which are kept completely isolated from the training and validation sets. All degraded data samples were randomly selected from the predefined degradation types and intensity levels specified in Appendix A.
@@ 5.2 Watermark Performance
Baseline Methods. We select seven baseline methods: three officially used by SD, namely DwtDct [old_wm-digital_wm_and_stega] and DwtDctSvd [old_wm-digital_wm_and_stega]; two post-processing-based methods RivaGAN [postprocessing_watermarking-RivaGAN] and EditGuard [wm_and_tl-Editguard], a fine tuning-based method called Stable Signature [fine_tuning_based_watermarking-StableSignature], a latent representation-based method called Tree-Ring [latent_representation_based_watermarking-Treering], and GS [gaussianshading].
Evaluation Metrics. To measure the performance of watermarking methods, we calculated the bit accuracy (Bit Acc). To measure the bias in model performance, we computed the CLIP-Score [CLIP] for 10 batches of watermarked images and performed a t-test on the mean CLIP-Score compared to that of watermark-free images. In prior works, the incorporation of watermark embedding modules inevitably results in a decline in model performance, therefore typically evaluated using the Peak Signal-to-Noise Ratio (PSNR) and Fréchet Inception Distance (FID) [metrics-FID], which are unnecessary to our method. To simulate detecting and tracing scenarios, we also calculated the detecting true positive rate (D-TPR) and the tracing true positive rate (T-TPR), by fixing the false positive rate (FPR) and the number of users at $10^{-6}$ and $10^{6}$, separately [gaussianshading].